# Investigating Solid-Fluid Phase Coexistence in DC Plasma Bilayer Crystals: The Role of Particle Pairing and Mode Coupling

### v1

@@ IV Conclusion
In this work, we explored the mechanisms driving solid–fluid phase coexistence and melting in a bilayer dusty plasma crystal in a DC discharge by combining experimental phonon spectra, particle-level dynamics, and a new quantitative measure of non-reciprocity. Our aim was to investigate the interplay between Mode-Coupling Instability (MCI), particle pairing, and structural disorder under varying confinement conditions.
Phonon spectral analysis revealed MCI-like hotspots near 19 Hz in the acoustic modes, while vertical oscillations in the optical modes occurred in the 30–39 Hz range well above the classical MCI resonance. These results confirm that instability in bilayers differs fundamentally from monolayer MCI, being strongly influenced by interlayer coupling and structural asymmetry. The Schweigert mechanism is also not applicable here: we observed clear vertical oscillations in both layers, growing with decreasing confinement bias, even at constant neutral pressure. This indicates that melting is governed by confinement-driven structural changes rather than a damping threshold.
@@
Together, these findings point to a hybrid mechanism for melting in bilayer dusty plasma crystals, where vertical confinement, wakefield-mediated non-reciprocal interactions, dynamic particle pairing, and local structural fluctuations act cooperatively to destabilize the crystalline order. This establishes dynamic particle pairing amplified by non-reciprocal energy injection as the key driver of the phase transition. The framework developed here, which combines phonon spectral analysis with direct quantification of non-reciprocity, provides a general approach for studying melting, energy transport, and defect dynamics in strongly coupled non-equilibrium systems.

### v3

@@ IV Conclusion
In this work, we explored the mechanisms driving solid-fluid phase coexistence and melting in a bilayer dusty plasma crystal in a DC discharge by combining experimental phonon spectra, particle-level dynamics, and a new quantitative measure of non-reciprocity. Our aim was to investigate the interplay between Mode-Coupling Instability (MCI), particle pairing, and structural disorder under varying confinement conditions.
Phonon spectral analysis revealed MCI-like hotspots near 19 Hz in the acoustic modes, while vertical oscillations in the optical modes occurred in the 30-39 Hz range well above the classical MCI resonance. These results confirm that instability in bilayers differs fundamentally from monolayer MCI, being strongly influenced by interlayer coupling and structural asymmetry. The Schweigert mechanism is also not applicable here: we observed clear vertical oscillations in both layers, growing with decreasing confinement bias, even at constant neutral pressure. This indicates that melting is governed by confinement-driven structural changes rather than a damping threshold.
@@
Together, these findings point to a hybrid mechanism for melting in bilayer dusty plasma crystals, where vertical confinement, wakefield-mediated non-reciprocal interactions, dynamic particle pairing, and local structural fluctuations act cooperatively to destabilize the crystalline order. These results indicate that dynamic particle pairing amplified by non-reciprocal energy injection is strongly correlated with the phase transition, and accompanies confinement-driven structural and fluctuation effects. The framework developed here, which combines phonon spectral analysis with direct quantification of non-reciprocity, provides a general approach for studying melting, energy transport, and defect dynamics in strongly coupled non-equilibrium systems.
V APPENDIX
Figure 11: Representative side view data ROI for 140V from which layer profiles were calculated.
Figure 12: Average vertical intensity profiles corresponding to the ROI in Fig. 11.
Figure 13: Vertical oscillations of the layers as seen from the side view. We can clearly see the layer interaction increase and the layer move closer as the ring bias voltage is reduced.
Analysis Methods
Each frame of the side-view video (stored as a TIFF stack) was analyzed using ImageJ. For the layer-resolved vertical dynamics shown in Figs. 13 and 13, the Region Of Interest (ROI) in Fig. 11 was used, and the analysis was performed directly on the image intensity profiles. For each frame, a vertical intensity profile was extracted using the Profile Plot function in ImageJ by averaging the pixel intensity across the horizontal extent of the ROI. The resulting one-dimensional intensity profiles exhibit distinct peaks corresponding to the vertically separated dust layers.
The peak positions were identified for each frame, providing the instantaneous vertical locations of the layers. These peak positions were then tracked in time to obtain the vertical trajectories shown in Fig. 13. The profiles shown in Fig. 13 correspond to horizontally-averaged vertical intensity profiles for each bias voltage at a fixed frame. Using the same ROI for all figures ensures that the inter-layer separation and layer-resolved oscillations are extracted from identical spatial regions of the dust cloud.
At a confinement ring bias voltage of 140 V, two well-resolved intensity peaks were consistently observed, indicating a bilayer configuration with a vertical separation large compared to the characteristic screening length, and therefore weak interlayer coupling.
At 120 V, the intensity profiles frequently exhibited three distinct peaks, reflecting a splitting of the upper layer into two vertically separated sublayers. The intermediate sublayer appeared within a vertical distance comparable to the effective screening length, resulting in enhanced interaction with both the upper and lower layers, while the direct interaction between the original top and bottom layers remained comparatively weak due to their larger separation.
At 103 V, all three intensity peaks shifted closer together, and the separations between adjacent layers decreased substantially. In this regime, the reduced vertical spacing indicates strong interlayer coupling among all layers, consistent with enhanced collective dynamics observed in both side-view and top-view measurements.